Solve the following system of linear equations for π‘₯, 𝑦, 𝑧, 𝑑: βˆ’π‘¦+3𝑧+2π‘₯+4𝑑 = 9

π‘₯βˆ’2𝑧+7𝑑 =11 π‘§βˆ’3𝑦+3π‘₯+5𝑑 = 8 𝑦 + 2π‘₯ + 4𝑧 + 4𝑑 = 10


1
Expert's answer
2021-10-29T03:01:54-0400
"\begin{matrix}
 2x-y+3z+4t=9 \\
 x-2z+7t=11\\
3x-3y+z+5t=8\\
2x+y+4z+4t=10
\end{matrix}"

Augmented matrix


"\begin{pmatrix}
 2 & -1 & 3 & 4 & & 9 \\
 1 & 0 & -2 & 7 & & 11 \\
 3 & -3 & 1 & 5 & & 8 \\
 2 & 1 & 4 & 4 & & 10 \\
\end{pmatrix}"

"R_2=R_2-R_1\/2"


"\begin{pmatrix}
 2 & -1 & 3 & 4 & & 9 \\
 0 & 1/2 & -7/2 & 5 & & 13/2 \\
 3 & -3 & 1 & 5 & & 8 \\
 2 & 1 & 4 & 4 & & 10 \\
\end{pmatrix}"

"R_3=R_3-3R_1\/2"


"\begin{pmatrix}
 2 & -1 & 3 & 4 & & 9 \\
 0 & 1/2 & -7/2 & 5 & & 13/2 \\
 0 & -3/2 & -7/2 & -1 & & -11/2 \\
 2 & 1 & 4 & 4 & & 10 \\
\end{pmatrix}"

"R_4=R_4-R_1"


"\begin{pmatrix}
 2 & -1 & 3 & 4 & & 9 \\
 0 & 1/2 & -7/2 & 5 & & 13/2 \\
 0 & -3/2 & -7/2 & -1 & & -11/2 \\
 0 & 2 & 1 & 0 & & 1 \\
\end{pmatrix}"

"R_3=R_3+3R_2"


"\begin{pmatrix}
 2 & -1 & 3 & 4 & & 9 \\
 0 & 1/2 & -7/2 & 5 & & 13/2 \\
 0 & 0 & -14 & 14 & & 14 \\
 0 & 2 & 1 & 0 & & 1 \\
\end{pmatrix}"

"R_4=R_4-4R_2"


"\begin{pmatrix}
 2 & -1 & 3 & 4 & & 9 \\
 0 & 1/2 & -7/2 & 5 & & 13/2 \\
 0 & 0 & -14 & 14 & & 14 \\
 0 & 0 & 15 & -20 & & -25 \\
\end{pmatrix}"

"R_4=R_4+15R_3\/14"


"\begin{pmatrix}
 2 & -1 & 3 & 4 & & 9 \\
 0 & 1/2 & -7/2 & 5 & & 13/2 \\
 0 & 0 & -14 & 14 & & 14 \\
 0 & 0 & 0 & -5 & & -10 \\
\end{pmatrix}"

"R_4=R_4\/(-5)"


"\begin{pmatrix}
 2 & -1 & 3 & 4 & & 9 \\
 0 & 1/2 & -7/2 & 5 & & 13/2 \\
 0 & 0 & -14 & 14 & & 14 \\
 0 & 0 & 0 & 1 & & 2 \\
\end{pmatrix}"

"R_3=R_3-14R_4"


"\begin{pmatrix}
 2 & -1 & 3 & 4 & & 9 \\
 0 & 1/2 & -7/2 & 5 & & 13/2 \\
 0 & 0 & -14 & 0 & & -14 \\
 0 & 0 & 0 & 1 & & 2 \\
\end{pmatrix}"

"R_3=R_3\/(-14)"


"\begin{pmatrix}
 2 & -1 & 3 & 4 & & 9 \\
 0 & 1/2 & -7/2 & 5 & & 13/2 \\
 0 & 0 & 1 & 0 & & 1 \\
 0 & 0 & 0 & 1 & & 2 \\
\end{pmatrix}"

"R_2=R_2-5R_4"


"\begin{pmatrix}
 2 & -1 & 3 & 4 & & 9 \\
 0 & 1/2 & -7/2 & 0 & & -7/2 \\
 0 & 0 & 1 & 0 & & 1 \\
 0 & 0 & 0 & 1 & & 2 \\
\end{pmatrix}"

"R_2=R_2+7R_3\/2"


"\begin{pmatrix}
 2 & -1 & 3 & 4 & & 9 \\
 0 & 1/2 & 0 & 0 & & 0 \\
 0 & 0 & 1 & 0 & & 1 \\
 0 & 0 & 0 & 1 & & 2 \\
\end{pmatrix}"

"R_2=2R_2"


"\begin{pmatrix}
 2 & -1 & 3 & 4 & & 9 \\
 0 & 1 & 0 & 0 & & 0 \\
 0 & 0 & 1 & 0 & & 1 \\
 0 & 0 & 0 & 1 & & 2 \\
\end{pmatrix}"

"R_1=R_1-4R_4"


"\begin{pmatrix}
 2 & -1 & 3 & 0 & & 1 \\
 0 & 1 & 0 & 0 & & 0 \\
 0 & 0 & 1 & 0 & & 1 \\
 0 & 0 & 0 & 1 & & 2 \\
\end{pmatrix}"

"R_1=R_1-3R_3"


"\begin{pmatrix}
 2 & -1 & 0 & 0 & & -2 \\
 0 & 1 & 0 & 0 & & 0 \\
 0 & 0 & 1 & 0 & & 1 \\
 0 & 0 & 0 & 1 & & 2 \\
\end{pmatrix}"

"R_1=R_1+R_2"


"\begin{pmatrix}
 2 & 0 & 0 & 0 & & -2 \\
 0 & 1 & 0 & 0 & & 0 \\
 0 & 0 & 1 & 0 & & 1 \\
 0 & 0 & 0 & 1 & & 2 \\
\end{pmatrix}"

"R_1=R_1\/2"


"\begin{pmatrix}
 1 & 0 & 0 & 0 & & -1 \\
 0 & 1 & 0 & 0 & & 0 \\
 0 & 0 & 1 & 0 & & 1 \\
 0 & 0 & 0 & 1 & & 2 \\
\end{pmatrix}"

The solution is


"x=-1, y=0, z=1, t=2"

"(x,y,z,t)=(-1,0,1,2)"